# Fractional diffusion as the limit of a short range potential Rayleigh gas

### v1

@@ 5.2. Estimate the loss term of the collision operator
In the remaining subsection, we are using [MMM11, Theorem 3.2] which we state it here for the convenience of the reader. Before stating the theorem, we appose here the key assumptions. That is,
@@
$\nu(v)F(v)=K(F)(v)=\int_{\mathds{R}^{3}}\sigma(v,v^{\prime})F(v^{\prime})\ \mathrm{d}v^{\prime},$
@@
Assumption $(A_{2})$: We set $G:=(1+\nu)F$, then $K(F)=\nu F$ is equivalent to $G=T(G)$, with $T:L^{1}(\mathds{R}^{3})\to L^{1}(\mathds{R}^{3})$ defined by
$T(G)(v)=\int_{\mathds{R}^{3}}\frac{1+\nu(v)}{1+\nu(v^{\prime})}\frac{\sigma(v,v^{\prime})}{\nu(v)}G(v^{\prime})\,\mathrm{d}v^{\prime}.$
By applying Krein-Rutman’s theorem, there exists a radial positive and unimodal equilibrium distribution $F$, see [DGP00, Section A], it is the unique normalised positive equilibrium.
Assumption $(B_{1})$: We know that the function $F$ is an equilibrium distribution of the linear Boltzmann collision operator (5.1), i.e. $\nu(v_{1})F(v_{1})=K(F)(v_{1})$. Additionally, by the computations in subsections 5.1 and 5.2 above we have that the behaviour of $F$ for large velocities is
$|v_{1}|^{q}F(v_{1})=|v_{1}|^{q}\frac{K(F)(v_{1})}{\nu(v_{1})}\to\frac{4C_{2}C_{S}}{(\frac{\pi}{2}-\sigma)^{2}}\quad as\ |v_{1}|\to\infty.$ (5.6)
@@
$|v_{1}|^{3+\alpha}F(v_{1})=|v_{1}|^{3+\alpha}\frac{K(F)(v_{1})}{\nu(v_{1})}\to\kappa_{0}\quad as\ |v_{1}|\to\infty$ (5.7)
By equating (5.6) and (5.7), we take that $\alpha=q-3>0$, which implies $q>3$ (actually, that $q=\alpha+3$, $\alpha>0$) which is true since $g_{0}\in L^{1}(\mathds{R}^{3})$ and also that $\kappa_{0}=\frac{4C_{2}C_{S}}{(\frac{\pi}{2}-\sigma)^{2}}$.
Assumption $(B_{2})$: By the behaviour of $\nu(v_{1})$ for large velocities $v_{1}$, in (5.5), choosing $\beta=\frac{n-5}{n-1}<1$, since $n>1$ and the positive constant $\nu_{0}=\tilde{C}_{K}$. Furthermore,
$\beta<\min\{\alpha;2-\alpha\}=\min\{q-3;5-q\}=5-q$
$\Leftrightarrow$ $\beta=\frac{n-5}{n-1}<5-q\Leftrightarrow q<\frac{4n}{n-1},\ \textrm{since}\ n>1.$
@@
$\nu(v)\int_{\mathds{R}^{3}}\frac{F(v^{\prime})}{b(v,v^{\prime})}\,\mathrm{d}v^{\prime}$ $=\nu(v)F(v)\int_{\mathds{R}^{3}}\frac{F(v^{\prime})}{\int_{\mathbb{S}^{2}}g_{0}(v)B(v-v^{\prime},\omega)\,\mathrm{d}\omega}\,\mathrm{d}v^{\prime}$
$=\frac{K(F)(v)}{g_{0}(v)}\int_{\mathds{R}^{3}}\frac{F(v^{\prime})}{\int_{\mathbb{S}^{2}}B(v-v^{\prime},\omega)\,\mathrm{d}\omega}\,\mathrm{d}v^{\prime},$
@@
$\int_{\mathbb{S}^{2}}B(v-v^{\prime},\omega)\,\mathrm{d}\omega$ $\geq\frac{\pi}{2}\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}(\frac{\pi}{2}-\sigma)^{2}|v-v^{\prime}|^{\frac{n-5}{n-1}}.$
@@
$\nu(v)\int_{\mathds{R}^{3}}\frac{F(v^{\prime})}{b(v,v^{\prime})}\,\mathrm{d}v^{\prime}$ $\leq 8\pi C_{2}C_{S}\frac{|v|^{-q+\frac{n-5}{n-1}}+o(|v|^{-q+\frac{n-5}{n-1}})}{|v|^{-q}}(|v|^{-\frac{n-5}{n-1}}+o(|v|^{-\frac{n-5}{n-1}}))$
@@
We are now ready to state the theorem which is the following.
@@
$\gamma:=\frac{\alpha-\beta}{1-\beta},\quad and\quad\theta(\varepsilon):=\ell(\epsilon^{-\frac{1}{1-\beta}})\epsilon^{\gamma}.$
@@
$\theta(\epsilon)\partial_{\tau}f+\epsilon v\cdot\nabla_{x}f=Q(f),$ (5.8)
@@
$\tau=\frac{\theta(\epsilon)}{\epsilon}t\quad with\quad c=\frac{1}{\epsilon}.$
@@
$c=\frac{1}{\epsilon}\quad\mathrm{and}\quad\tilde{\beta}=\frac{\epsilon}{\theta(\epsilon)},$
@@
$|(\hat{f}^{N}(t,x,v)-f(t,x,v))\cdot\phi(t,x,v)|\to 0,$
as $N\to\infty$, which by the Boltzmann-Grad limit is equivalent to $\varepsilon\to 0$. Now by the Theorem 5.3 we know that the rescaled solution of the linear Boltzmann equation $f^{\epsilon}(\theta(\epsilon)t,\epsilon x,v)$ converges weakly in $L^{\infty}(0,T;L^{2}_{F^{-1}}(\mathds{R}^{3}\times\mathds{R}^{3}))$ to a function $\rho(t,x)F(v)$, where $\rho(t,x)$ is the unique solution of the fractional diffusion equation (1.7) and the function $F$ is an equilibrium distribution of the linear Boltzmann operator (5.1), i.e. $Q(F)=0$. That is to say, for any test function $\psi\in L^{1}(0,T;L^{2}_{F^{-1}}(\mathds{R}^{3}\times\mathds{R}^{3}))$
$|(f^{\epsilon}(\theta(\epsilon)t,\epsilon x,v)-\rho(t,x)F(v))\cdot\psi(t,x,v)|\to 0,$
@@
$|(\hat{f}^{N}(\tilde{\beta}\tau,x,v)-\rho(\tau,x)F(v))\cdot\tilde{\varphi}(\tau,x,v)|\to 0,$

### v2

@@ 5.3. Existence of Equilibrium
In the remaining subsection, we are using [MMM11, Theorem 3.2] which we state it here for the convenience of the reader. Before stating the theorem, we list here the key assumptions.
@@
[…]
@@
Before stating the next assumption, we recall the definition of slowly varying functions. A measurable function $\ell:(0,\infty)\to(0,\infty)$ is called slowly varying (at infinity) if for all $a>0$, $\lim_{x\to\infty}{\frac{\ell(ax)}{\ell(x)}}=1.$
@@
Assumption $(A_{2})$: This follows directly from Proposition 5.6.
Assumption $(B_{1})$: We know that the function $F$ is an equilibrium distribution of the linear Boltzmann collision operator (5.2), i.e. $\nu(v_{1})F(v_{1})=K(F)(v_{1})$. Additionally, by the computations in subsections 5.2 and 5.1 above we have that the behaviour of $F$ for large velocities is
$|v_{1}|^{q}F(v_{1})=|v_{1}|^{q}\frac{K(F)(v_{1})}{\nu(v_{1})}\to\frac{16C_{2}C_{S}}{\pi^{2}}\quad as\ |v_{1}|\to\infty.$ (5.16)
@@
$|v_{1}|^{3+\alpha}F(v_{1})=|v_{1}|^{3+\alpha}\frac{K(F)(v_{1})}{\nu(v_{1})}\to\kappa_{0}\quad as\ |v_{1}|\to\infty.$ (5.17)
By equating (5.16) and (5.17), we take that $\alpha=q-3>0$, which implies $q>3$ (actually, that $q=\alpha+3$, $\alpha>0$) which is true since $g_{0}\in L^{1}(\mathds{R}^{3})$ and also that $\kappa_{0}=\frac{16C_{2}C_{S}}{\pi^{2}}$.
Assumption $(B_{2})$: By the behaviour of $\nu(v_{1})$ for large velocities $v_{1}$, in (5.7), choosing $\beta=\frac{n-5}{n-1}<1$, since $n>1$ and the positive constant $\nu_{0}=\tilde{C}$. Furthermore,
[…]
@@
[…]
$=\frac{K(F)(v)}{g_{0}(v)}\int_{\mathds{R}^{3}}\frac{F(v^{\prime})}{\int_{\mathbb{S}^{2}}\mathcal{B}(v-v^{\prime},\omega)\,\mathrm{d}\omega}\,\mathrm{d}v^{\prime},$
@@
$\int_{\mathbb{S}^{2}}\mathcal{B}(v-v^{\prime},\omega)\,\mathrm{d}\omega$ $\geq\frac{\pi^{3}}{8}\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}|v-v^{\prime}|^{\frac{n-5}{n-1}}.$
@@
[…]
@@
We are now ready to state the theorem of [MMM11] which is the following.
@@
$\gamma:=\frac{\alpha-\beta}{1-\beta},\quad\text{and}\quad\theta(\epsilon):=\ell(\epsilon^{-\frac{1}{1-\beta}})\epsilon^{\gamma}.$
@@
In our case, we have chosen $\ell=1$, so that $\theta(\epsilon)$ becomes explicitly $\theta(\epsilon)=\epsilon^{\gamma}$. Furthermore, the admissible choices for $\alpha$ and $\beta$ are
$\alpha=q-3,\ \text{for}\ q\in(4,\frac{4n}{n-1})\ \text{and}\ \beta=\frac{n-5}{n-1},\ \text{for}\ n\in(3,5].$
Therefore,
$\gamma=\frac{(q-4)(n-1)}{4}+1,\ \text{for}\ n\in(3,5],\ q\in(4,\frac{4n}{n-1}).$
@@
$\epsilon^{\gamma}\partial_{\tau}f+\epsilon v\cdot\nabla_{x}f=Q(f),$ (5.18)
@@
$\tau=\epsilon^{\gamma-1}t\quad\text{with}\quad c=\frac{1}{\epsilon}.$
@@
$c=\frac{1}{\epsilon}\quad\mathrm{and}\quad\tilde{\beta}=\epsilon^{1-\gamma},$
@@
$\int_{0}^{T}\int_{\mathds{R}^{3}\times\mathds{R}^{3}}|(\hat{f}^{N}(t,x,v)-f(t,x,v))\cdot\phi(t,x,v)|\mathrm{d}x\mathrm{d}v\mathrm{d}t\to 0,$
as $N\to\infty$, which by the Boltzmann-Grad limit is equivalent to $\varepsilon\to 0$. Now by the Theorem 5.9 we know that the rescaled solution of the linear Boltzmann equation $f^{\epsilon}(\epsilon^{\gamma}t,\epsilon x,v)$ converges weakly in $L^{\infty}(0,T;L^{2}_{F^{-1}}(\mathds{R}^{3}\times\mathds{R}^{3}))$ to a function $\rho(t,x)F(v)$, where $\rho(t,x)$ is the unique solution of the fractional diffusion equation (1.6) and the function $F$ is an equilibrium distribution of the linear Boltzmann operator (5.2), i.e. $Q(F)=0$. That is to say, for any test function $\psi\in L^{1}(0,T;L^{2}_{F^{-1}}(\mathds{R}^{3}\times\mathds{R}^{3}))$
$\int_{0}^{T}\int_{\mathds{R}^{3}\times\mathds{R}^{3}}|(f^{\epsilon}(\epsilon^{\gamma}t,\epsilon x,v)-\rho(t,x)F(v))\cdot\psi(t,x,v)|\mathrm{d}x\mathrm{d}v\mathrm{d}t\to 0,$
@@
$\int_{0}^{T}\int_{\mathds{R}^{3}\times\mathds{R}^{3}}|(\hat{f}^{N}(\epsilon^{1-\gamma}\tau,x,v)-\rho(\tau,x)F(v))\cdot\tilde{\varphi}(\tau,x,v)|\mathrm{d}x\mathrm{d}v\mathrm{d}\tau\to 0,$
@@ Acknowledgements
Data Availibility
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Conflict of interest
The authors have no relevant financial or non-financial interests to disclose.